Please solve it step by step with clear handwriting!.
Example Problem (5.2) A 75 kg skier starts down a 50-m-high, 10° slope on frictionless skis. What is his speed at the bottom? DOLL E R DF G H Shift Alt Car Ctri
Answer #1

The speed at the bottom will be

v = \sqrt{2gh} = \sqrt{2\times9.8\times 50} = 31.30m/s